# Reflectivity Is All You Need!: Advancing LiDAR Semantic Segmentation

### v1

@@ III-B3 Urban Environment
For urban environment testing, we utilize the SemanticKITTI dataset, a multi-modal dataset designed for urban environments, which features annotated LiDAR scans from the Velodyne HDL-64E. The performance on SemanticKITTI is detailed in Table II, highlighting the efficacy of different input configurations. The configuration denoted as $rxyzirn$ achieved the highest mean Intersection over Union (mIoU) at $54.8\%$, followed closely by $rxyzn$ at $54.2\%$, and the baseline configuration $rxyzi$, which also demonstrated competitive performance, achieving a $54.0\%$ mIoU.
@@ III-C Qualitative Results
Qualitative analysis was performed on Rellis-3D, with SalsaNext input configurations $rxyzi$ and $rxyzn$ as shown in Figure 6. Figure 6 compares predictions with ground truth represented in spherical projection and camera projection of segmented object points. (Note: SalsaNext doesn’t use images for training, and the figure is only for visualization purposes.) The figure shows that the predictions from $rxyzn$ provide better boundaries for puddles and dirt than predictions from $rxyzi$ and even ground truth. Moreover, the ground truth labels dirt surrounding the puddle region as a puddle, and many more instances of incorrect labeling can be found in the Rellis-3D dataset. This could be attributed to the overall low performance of SalsaNext across all configurations.
@@
III-D Runtime Evaluation
@@ IV Inference
The experiments conducted on the Rellis-3D and SemanticKITTI datasets indicate that utilizing reflectivity rather than intensity yields improved segmentation results in both off-road and on-road contexts. The notable enhancement in accuracy observed in off-road scenes can be attributed to the reflectivity parameter’s ability to distinguish between classes. This is particularly crucial as objects within the same class maintain consistent color and texture, while variations exist across different classes. Given the unstructured nature of off-road environments and the presence of loose boundaries, relying solely on geometry-based features resulted in lower performance.
In urban settings, where boundaries are clearly delineated across scenes, geometric features play a more significant role in learning compared to reflectivity/intensity parameters. Additionally, variations in intensity values between objects, such as different intensity values for cars based on surface color and texture, can also impact segmentation results. These factors may contribute to the limited improvement observed when using reflectivity in SemanticKITTI experiments. Nevertheless, it is noteworthy that incorporating reflectivity consistently led to enhanced segmentation results compared to intensity across all experiments.
After conducting cross-sensor experiments on the Rellis-3D dataset, it is clear that transferring a model from one sensor to another, even within the same scene, produces unsatisfactory outcomes. Although LiDAR intensity calibration has been effective in enhancing results, other factors such as the number of rings, ray pattern, and field of view have notable impacts on cross-domain capability.

### v2

@@ III-B3 Urban Environment
For urban environment testing, we utilize the SemanticKITTI dataset, a multi-modal dataset designed for urban environments, which features annotated LiDAR scans from Velodyne HDL-64E. The performance on SemanticKITTI is detailed in Table III, highlighting the efficacy of different input configurations. The configuration indicated as $rxyzirn$ achieved the highest mIoU at $54.8\%$, followed closely by $rxyzn$ at $54.2\%$, and the baseline configuration $rxyzi$, which also demonstrated competitive performance, achieving a $54.0\%$ mIoU.
@@ III-C Qualitative Results
Qualitative analysis was performed on Rellis-3d, with SalsaNext input configurations $rxyzi$ and $rxyzn$ as shown in Figure 5. Figure 5 compares the predictions with the ground truth represented in spherical projection and camera projection of segmented object points. (Note: SalsaNext doesn’t use images for training, and the figure is only for visualization purposes.) The figure shows that the predictions from $rxyzn$ provide better boundaries for puddles and dirt than the predictions of $rxyzi$ and even ground truth. In addition, the ground truth labels the dirt surrounding the puddle region as a puddle, and many more instances of incorrect labeling can be found in the Rellis-3d dataset. This could be attributed to the overall low performance of SalsaNext in all configurations.
@@
Figure 6: SAM inputs and corresponding annotated masks. The prompts are single point coordinate on the object.
III-D Reflectivity on Segment Anything
In this experiment, we utilized the Segment Anything Model (SAM), a foundation model designed for segmentation via zero-shot generalization using prompts. Specifically, we employed the ViT-H SAM checkpoint, which was trained on over 1 billion segmentation masks across 11 million RGB images. To adapt SAM for LiDAR data, we generated LiDAR images by spherically projecting range, intensity, and reflectivity values, using these as inputs in place of the traditional RGB channels.
As illustrated in Figure 6, SAM produced segmentation masks for the LiDAR inputs, which were subsequently annotated with corresponding labels. The objective of this experiment was to demonstrate the capability of SAM to generate segmentation masks for objects, highlighting the enhanced data representation provided by LiDAR scans through reflectivity. Given SAM’s widespread application in image annotation tasks [23], we present its use case in annotating LiDAR scans, a process that has historically been labor-intensive and time-consuming.
III-E Runtime Evaluation
@@ IV Inference
Experiments conducted on the Rellis-3D and SemanticKITTI datasets demonstrate that utilizing reflectivity instead of intensity leads to improved segmentation results in both off-road and on-road scenarios. The significant accuracy boost observed in off-road scenes can be attributed to reflectivity’s superior ability to discriminate between classes. This is crucial as objects within a class may share color and texture, while variations exist across classes. The unstructured nature of off-road environments with loose boundaries makes geometry-based features less reliable, leading to lower performance when used alone.
In contrast, in urban settings with well-defined boundaries, geometric features contribute more significantly to learning compared to reflectivity/intensity. Variations in intensity between objects of the same class (e.g., cars with different surface colors) can also affect segmentation. This may explain the limited improvement seen with reflectivity in SemanticKITTI. Nonetheless, incorporating reflectivity consistently improved segmentation compared to intensity across all experiments.
Cross-sensor experiments on Rellis-3D confirm that transferring a model between sensors, even within the same scene, yields poor results. While LiDAR intensity calibration helps, factors like ring count, ray pattern, and field of view significantly impact cross-domain capability.